Debbie's Cookies has a return on assets of 9.5 percent and a cost of equity of 13.9 percent. What is the pretax cost of debt if the debt–equity ratio is .73? Ignore taxes.
Answer #1

Debt-equity ratio=debt/equity

Hence debt=0.73 equity

Let equity be $x

Debt=$0.73x

Total=$1.73x

ROA=Respective costs*Respective weight

9.5=(x/1.73x*13.9)+(0.73x/1.73x*Cost of debt)

9.5=8.034682081+(0.73/1.73*Cost of debt)

Cost of debt=(9.5-8.034682081)*(1.73/0.73)

=3.47%(Approx).